A 10cm×10cm×10cm box contains 0.010 mol of nitrogen at 20∘C.

What is the rate of collisions (collisions/s) on one wall of the box?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Number of particles, N = 0.01 x 6.023 x 10^23 = 6.023 x 10^21

Mass, m = (Molecular mass of N)/6.023 x 10^23 = 4.65 x 10^-26 kg

Velocity, v_rms = sqrt(3 x 1.38 x 10^-23 x (273 + 20)/28 (1.67 x 10^-27)) = 511 m/s

Velocity, v = sqrt(V_rms^2/3) = 295 m/s

Rate of collision = (6.023 x 10^21 x 295)/(2 x 0.1) = 8.86 x10^25

Comment in case any doubt please rate my answer...